OpenStudy (anonymous):

A right triangle has one side of length 7 and a hypotenuse of length 10. How long is the third side of the triangle?

OpenStudy (anonymous):

Pythagorean Theorem: (leg)^2 + (leg)^2 = (hypotenuse)^2 7^2 + (leg)^2 = 10^2 49 + leg^2 = 100 leg^2 = 51 leg = sqrt (51)